OpenStudy (spring98):

Maya solved an equation incorrectly, as shown below: Step 1: 4x = 28 Step 2: x = 28 – 4 Step 3: x = 24 Which statement best explains why Step 2 is incorrect in Maya's solution? She did not add 4 to 28. She did not multiply 28 by 4. She did not divide 28 by 4. She did not multiply 28 four times.

Nnesha (nnesha):

4 times x = 28 to cancel oout 4 what would u do ? subtract add multiply or divide ?

OpenStudy (michele_laino):

Maya has to divide by 4, not subtract 4

OpenStudy (spring98):

divide 4 with what number?

OpenStudy (michele_laino):

4 is a factor not an addend so Maya has to divide both sides of her equation by 4, like this: \[\Large \frac{{4x}}{4} = \frac{{28}}{4}\]

OpenStudy (spring98):

the x would equal 7. so 4*7=28

OpenStudy (michele_laino):

no, the value of x, is: x=28/4=... since we have: